### v2

@@ 5.1 A Nonparametric Estimation Formulation for S-RAM
$\sum_{j\in\mathcal{I}^{(k)}}a_{ij}x_{j}^{k}\leq b_{i},\qquad\qquad\qquad\qquad\qquad\quad\forall k\in\mathcal{K},\forall i\in[m],$ (8)
@@
5.2 A Compact-sized Mixed Integer Convex Reformulation for Estimation of S-RAM
@@
$\delta_{j}^{k,l}-1\ \leq\ x_{j}^{l}-x_{j}^{k}\ \leq\ 1-\delta_{j}^{l,k},\qquad\qquad\qquad\qquad\quad\;\forall k,l\in\mathcal{K},\forall j\in\mathcal{I}^{(l)}\cap\mathcal{I}^{(k)},$
- $(\delta_{j}^{k,l}+\delta_{j}^{l,k})\ \leq\ x_{j}^{l}-x_{j}^{k}\ \leq\ \delta_{j}^{k,l}+\delta_{j}^{l,k},\ \qquad\qquad\qquad\;\,\forall k,l\in\mathcal{K},\forall j\in\mathcal{I}^{(l)}\cap\mathcal{I}^{(k)},$ (9)
@@
$x_{j}^{k}\ =\ x_{j}^{l}\quad\text{if}\quad\sum_{i\in[m]}{\lambda_{i}^{k}}_{\ast}a_{ij}=\sum_{i\in[m]}{\lambda_{i}^{l}}_{\ast}a_{ij},\qquad\quad\!\!\forall k,l\in\mathcal{K},\forall j\in\mathcal{I}^{(k)}\cap\mathcal{I}^{(l)},$ (10)
@@ 6 Numerical Experiments
In this section, we generate synthetic datasets to evaluate (i) the explanatory ability of S-RAM as captured by the representability in Step 1 and the best-fitting cost of S-RAM data in Step $1^{\prime}$, and (ii) the efficacy of predictions obtained in Step 2 of the workflow in Figure 1 for a wide range of network problems including longest path, shortest path, and assignment. All details on experiment implementation in this section are provided in Section B. We use a MacBook Pro Laptop with a 2GHz 4-core Intel Core i5 processor for all experiments.
@@
Representation and estimation under model misspecification. We evaluate the representational power of S-RAM on three combinatorial problems: longest path, shortest path, and assignment with observed data outside the S-RAM family. We generate directed acyclic graphs (DAGs) with 10 nodes for path problems and bipartite graphs with 4 and 6 nodes on each side for assignment problems. We vary the number of historical graphs across $K\in\{5,10,15,20,25,30\}$. Edges are included with probability drawn uniformly from $[0.4,0.6]$ while ensuring source-to-sink connectivity in path problems and at least one link per node in the assignment problem. On average, there are around 20 nodes in the DAGs and around 12 edges in the bipartite graphs.
@@
We generate synthetic instances outside the S-RAM framework by assigning edge costs in the form $v_{e}+\epsilon_{e}$, where $v_{e}$ is a deterministic component drawn uniformly from $[1,7]$, and $\epsilon_{e}$ is a Gaussian noise term. We consider three noise structures: independent, negatively correlated (correlation coefficients drawn from [0.9,0.5][-0.9,-0.5]), and positively correlated (from $[0.5,0.9]$), with variances sampled uniformly from $[9,25]$. For each graph, we solve the deterministic version of the problem 100 times under different noise realizations and average the solutions to construct the observed data.
For each instance, we verify S-RAM representability for it via (3) and, if violated, compute the average absolute deviation loss using (9). Then for each $K$ and correlation structure, we report the representable proportion and the average absolute deviation loss over 20 instances. As shown in Table 1, S-RAM performs well across all tested problems despite model misspecification, achieving high representability for $K\leq 15$ and low estimation error over all testing cases. The framework is robust across correlation structures, with better performance under independent and negatively correlated noise. Assignment yields the highest representability and lowest loss overall.
@@
Table 1: Representation and estimation across longest path, shortest path, and assignment problems with Gaussian instances.
Problem K independent gaussian gaussian with negative correlation gaussian with positive correlation
represent-
able
propotion
average
absolute
deviation
loss (se)
@@
represent-
able
propotion
average
absolute
deviation
loss (se)
@@
represent-
able
propotion
average
absolute
deviation
loss (se)
@@
Longest
Path
5 0.95 0.10 (0.10) 0.85 6.77 (4.60) 1 0.00 (0.00)
10 0.85 7.80 (5.73) 0.65 6.23 (2.29) 0.4 13.08 (5.79)
15 0.5 9.69 (4.19) 0.5 22.86 (7.19) 0.35 13.19 (5.48)
20 0.3 38.35 (14.34) 0.4 27.13 (11.23) 0.15 34.84 (10.36)
25 0.2 52.01 (14.42) 0.1 66.44 (20.72) 0.05 49.76 (10.61)
30 0.05 57.04 (12.82) 0.05 72.90 (17.47) 0 92.88 (20.97)
Shortest
Path
5 0.9 0.31 (0.26) 0.9 0.16 (0.12) 0.95 0.02 (0.02)
10 0.65 3.98 (2.25) 0.8 0.85 (0.83) 0.65 2.67 (1.38)
15 0.6 5.41 (3.18) 0.65 3.97 (3.13) 0.45 5.41 (2.90)
20 0.35 6.16 (2.63) 0.5 2.04 (0.88) 0.15 17.36 (6.80)
25 0.1 27.15 (7.41) 0.15 13.67 (3.11) 0 33.96 (6.61)
30 0.15 22.91 (6.99) 0.15 27.87 (6.81) 0.05 56.68 (15.24)
Assign-
ment
5 0.95 0.03 (0.02) 1 0.00 (0.00) 0.95 0.01 (0.00)
10 0.8 0.16 (0.12) 0.95 0.14 (0.14) 0.75 0.32 (0.29)
15 0.85 0.61 (0.37) 0.65 0.37 (0.27) 0.8 0.03 (0.02)
20 0.7 0.48 (0.35) 0.55 1.71 (0.88) 0.55 0.39 (0.23)
25 0.55 3.97 (2.14) 0.45 1.09 (0.60) 0.7 1.05 (0.65)
30 0.4 2.06 (0.85) 0.4 0.74 (0.38) 0.3 2.38 (1.61)
@@
Prediction accuracy with S-RAM data. We evaluate the predictive accuracy of nonparametric S-RAM on counterfactual tasks using data generated from parametric S-RAMs. For the path problems, we test whether S-RAM can predict the flow on a key edge that consistently appears in the optimal path when the graph changes to a new structure. For the assignment problem, we examine whether S-RAM can predict the total system welfare on a new graph with some nodes being unavailable.
Since S-RAM and MDM are equivalent, to retain interpretability in utility theory, we generate data from parametric MDMs. To assess the robustness of nonparametric S-RAM, we generate MDM data with different parametric assumptions on marginals. In path problems, marginals of noise follow nonidentical exponential distributions with rates from $[1,2]$. To ensure that the edge of interest is influential in the optimal solution, we set its deterministic cost to 8 in the longest path problem and to 1 in the shortest path problem. All other edge costs are sampled uniformly from $[1,7]$. We follow similar ways to generate DAGs for path problems. In assignment, marginals of noise follow nonidentical uniform distributions in the form of $\text{Uniform}(a,b)$ and $a$ being sampled uniformly from $[1,2]$ and $b$ being sampled uniformly from $[5,10]$. We generate the bipartite graphs by randomly assigning some nodes to be unavailable and removing all the associated edges. We evaluate the prediction performance of nonparametric S-RAM across collection sizes $K\in\{5,10,20\}$. For each value of $K$, we generate 20 test instances for each problem. For path problems, we compute the S-RAM worst- and best-case estimates for the flow on the selected edge. For assignment, predicting the worst-case welfare leads to a nonconvex mixed integer program, which is computationally challenging and might not be accurate when solving by existing solvers. We use the best-case estimate for the total welfare. The prediction results are reported in Figure 2.
The figure shows that the proposed nonparametric approach can accurately recover the true flow or total welfare, with prediction intervals that narrow as the number of historical graphs increases and the predicted total welfares that closely match the ground truths across all test instances. Overall, our experiments show that nonparametric S-RAM effectively captures the ground truth across diverse combinatorial structures.

### v3

@@ 5.1. A Nonparametric Estimation Formulation for s-RAM
(11) $\sum_{j\in\mathcal{I}^{(k)}}a_{ij}x_{j}^{k}\leq b_{i},\qquad\qquad\qquad\qquad\qquad\quad\forall k\in\mathcal{K},\forall i\in[m],$
@@
5.2. A Compact-sized Mixed Integer Convex Reformulation for Estimation of s-RAM
@@
$\delta_{j}^{k,l}-1\ \leq\ x_{j}^{l}-x_{j}^{k}\ \leq\ 1-\delta_{j}^{l,k},\qquad\qquad\qquad\qquad\;\;\,\forall k,l\in\mathcal{K},\forall j\in\mathcal{I}^{(l)}\cap\mathcal{I}^{(k)},$
(12) - $(\delta_{j}^{k,l}+\delta_{j}^{l,k})\ \leq\ x_{j}^{l}-x_{j}^{k}\ \leq\ \delta_{j}^{k,l}+\delta_{j}^{l,k},\ \qquad\qquad\qquad\,\forall k,l\in\mathcal{K},\forall j\in\mathcal{I}^{(l)}\cap\mathcal{I}^{(k)},$
@@
(13) $x_{j}^{k}\ =\ x_{j}^{l}\quad\text{if}\quad\sum_{i\in[m]}{\lambda_{i}^{k}}_{\ast}a_{ij}=\sum_{i\in[m]}{\lambda_{i}^{l}}_{\ast}a_{ij},\qquad\quad\!\!\forall k,l\in\mathcal{K},\forall j\in\mathcal{I}^{(k)}\cap\mathcal{I}^{(l)},$
@@ 6. Numerical Experiments
In this section, we generate synthetic datasets to evaluate (i) the explanatory ability of s-RAM as captured by the representability in Step 1 and the best-fitting cost of s-RAM data in Step $1^{\prime}$, and (ii) the efficacy of predictions obtained in Step 2 of the workflow in Figure 1 for a wide range of combinatorial problems including longest path, shortest path, linear assignment, and constrained subset selection problems. Missing details on experiment implementation in this section are provided in Section B. We use a MacBook Pro with a 12-core Apple M4 Pro chip and 24 GB RAM for all experiments.
6.1. Longest Path Problem
6.1.1. Explanatory ability of s-RAM in the longest path problem
We generate data from models that lie outside the s-RAM framework to assess the explanatory ability of s-RAM under model misspecification.
@@ 6.1.1. Explanatory ability of s-RAM in the longest path problem
Generation of historical $\mathcal{X}^{(k)}$ instances. We take the number of historically observed graph instances to vary across the range $K\in\{5,10,15,20\}$. We generate directed acyclic graphs (DAGs) with 10 nodes. Edges are included with probability drawn uniformly from $[0.4,0.6]$ while ensuring source-to-sink connectivity. On average, there are around 25 edges in the DAGs.
@@
Generation of historical $\boldsymbol{p}^{(k)}$ instances. We generate the respective synthetic $\boldsymbol{p}^{(k)}$ data values outside the s-RAM framework by assigning edge costs in the form $v_{e}+\epsilon_{e}$, where $v_{e}$ is a deterministic component drawn uniformly from $[1,7]$, and $\epsilon_{e}$ is a Gaussian noise term. We consider three noise structures: independent, negatively correlated (correlation coefficients drawn from [0.9,0.5][-0.9,-0.5]), and positively correlated (from $[0.5,0.9]$), with variances sampled uniformly from $[9,25]$. For each graph, we solve the deterministic version of the problem 100 times under different noise realizations and average the solutions to construct the observed data.
Representability test and estimation with $\boldsymbol{p}^{\mathcal{K}}$. For each instance, we first verify whether the observed data are s-RAM representable by solving (7). If an instance is not s-RAM representable, we quantify the degree of violation by computing the average absolute deviation loss. For the longest path problem, the resulting estimation problem corresponding to (12) with the chosen loss function can be written as
(14) $\begin{aligned} \min_{\boldsymbol{x},\boldsymbol{\lambda},\boldsymbol{\delta}\!\!\,}\quad&\quad\frac{1}{K}\sum_{k=1}^{K}\sum_{(i,j)\in\boldsymbol{E}^{k}}p_{ij}^{k}|x_{ij}^{k}-p_{ij}^{k}|\\ \text{s.t.}\quad\ -&\delta^{k,l}_{i,j}\ \leq\ \lambda_{j}^{k}-\lambda_{i}^{k}-\lambda_{j}^{l}+\lambda_{i}^{l}\ \leq\ 1-(1+\epsilon)\delta^{k,l}_{i,j},\quad\ \ \forall\,k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},\\ &\delta^{k,l}_{i,j}-1\ \leq\ x_{ij}^{l}-x_{ij}^{k}\ \leq\ 1-\delta^{l,k}_{i,j},\quad\qquad\quad\forall\,k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},\\ -&(\delta^{k,l}_{i,j}+\delta^{l,k}_{i,j})\ \leq\ x_{ij}^{l}-x_{ij}^{k}\ \leq\ \delta^{k,l}_{i,j}+\delta^{l,k}_{i,j},\ \quad\forall\,k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},\\ &\sum_{j:(i,j)\in\boldsymbol{E}^{k}}x_{ij}-\sum_{j:(j,i)\in\boldsymbol{E}^{k}}x_{ji}=\begin{cases}1,\quad&i=s,\\ -1,\quad&i=t,\\ 0,\quad&i\in\boldsymbol{V}^{k},\end{cases}\forall k\in\mathcal{K},\\ &x_{ij}^{k}\geq 0,\ \forall\,k\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k},\\ &\delta^{k,l}_{i,j}\,\in\,\{0,1\},\ \forall\,k,l\in\mathcal{K},\ \forall(i,j)\in\boldsymbol{E}^{k}\cap\boldsymbol{E}^{l},\end{aligned}$
where $0<\epsilon<1/(2|\boldsymbol{V}|K)$ and $\boldsymbol{V}$ is the collection of all possible nodes. For each value of $K$ and each correlation structure, we report in Table 2 the fraction of s-RAM-representable instances and the corresponding average absolute deviation loss, computed over 20 independently generated instances. Because the data are generated randomly, enforcing exact representability can be overly restrictive: instances with vanishingly small loss may be classified as non-representable due to numerical or sampling noise. Accordingly, we treat an instance as s-RAM representable if its average absolute deviation loss is below $10^{-6}$, and report the representable fraction under this criterion.
Representation and estimation results in the longest path problem. Table 2 shows that s-RAM performs well across all tested scenarios despite model misspecification, achieving high representability for $K\leq 15$. For instance, when $K=5$, more than $85\%$ of the instances are s-RAM representable across all correlation structures, and the representable fraction remains above $60\%$ even when $K=15$. Moreover, the associated estimation error remains small: the average absolute deviation loss is consistently on the order of $10^{-3}$ or smaller, even in settings where exact representability is violated. Naturally, the representability decreases with $K$ due to the increasing number of feasibility constraints in (6) induced by additional historical observations. Overall, we observe the framework to be robust across different correlation structures.
@@
Table 2. Representation and estimation in the longest path problem with Gaussian instances
K independent gaussian gaussian with -ve correlation gaussian with +ve correlation
fraction
of s-RAM
representable
instances
average absolute
deviation loss (se)
@@
fraction
of s-RAM
representable
instances
average absolute
deviation loss (se)
@@
fraction
of s-RAM
representable
instances
average absolute
deviation loss (se)
@@
5 0.95 0.02 (0.02) 0.85 1.35 (0.92) 1.00 0.00 (0.00)
10 0.85 0.78 (0.57) 0.70 0.62 (0.23) 0.45 1.31 (0.58)
15 0.70 0.65 (0.28) 0.60 1.52 (0.48) 0.55 0.88 (0.37)
20 0.45 1.92 (0.72) 0.45 1.36 (0.56) 0.30 1.74 (0.52)
6.1.2. Predictive ability of s-RAM in the longest path problem
@@
We evaluate the predictive accuracy of nonparametric s-RAM on counterfactual tasks using data generated from parametric s-RAMs. For the longest path, we test whether s-RAM can predict the probability of an edge being included in the critical path when the graph changes to a new structure.
Generation of training instances $\boldsymbol{p}^{(k)}$ and ground truth for the longest path problem. Since s-RAM and MDM are equivalent, to retain interpretability in utility theory, we generate data from parametric MDMs. In the longest path, marginals of noise follow nonidentical exponential distributions with rates sampled from $[1,2]$. To ensure that the edge of interest is influential in the optimal solution, we set its deterministic cost to $8$. All other edge costs are sampled independently and uniformly from $[1,7]$. We generate directed acyclic graphs following the same procedure as in the explanatory ability test above. Then, for each instance $k$, we use formulation (18) to generate training instances $\boldsymbol{p}^{(k)}.$ We evaluate the prediction performance of the nonparametric s-RAM across collection sizes $K\in\{5,10,20\}$. For each value of $K$, we generate $20$ training instances. Given a new graph $\boldsymbol{G}^{\text{new}}=(s\cup t\cup\boldsymbol{V}^{\text{new}},\boldsymbol{E}^{\text{new}})$. Let $\boldsymbol{x}^{\text{true}}$ denote the resulting vector of edge inclusion probabilities for $\boldsymbol{G}^{\text{new}}.$ Then we use the same parameter values as in the training instances together formulation (18) to generate $\boldsymbol{x}^{\text{true}}$. Let $(i^{*},j^{*})$ be the edge of interest. The corresponding ground-truth quantity is then $x^{\text{true}}_{i^{*}j^{*}}$.
Prediction with nonparametric s-RAM in the longest path. For the longest path problem, we compute the s-RAM worst- and best-case estimates for the probability that the selected edge $(i^{*},j^{*})$ is included in the critical path. Since the observed data $\boldsymbol{p}^{\mathcal{K}}$ are representable by s-RAM, given a new graph $\boldsymbol{G}^{\text{new}}=(s\cup t\cup\boldsymbol{V}^{\text{new}},\boldsymbol{E}^{\text{new}})$, Proposition 4.2 implies that the worst-case estimate $\underline{f}_{\text{new}}$ can be obtained by solving the following mixed-integer linear program, with $0<\epsilon<1/(2|\boldsymbol{V}|K)$:
(15) $\min_{\boldsymbol{x},\boldsymbol{\lambda},\boldsymbol{\delta}\!\!\,}\quad$ $\quad x_{i^{*}j^{*}}$
$x_{ij}\geq 0,\ \forall k\in\mathcal{K},(i,j)\in\boldsymbol{E}^{k},$
Likewise, the best-case estimate for the objective value $\bar{f}_{\text{new}}$ equals the optimal value obtained by maximizing the same objective over the constraints in the above mixed integer linear program.
The prediction results reported in Figure 2 show that the proposed nonparametric approach accurately recovers the ground truth. The estimated probability of the selected edge concentrates around the true probability, and the resulting prediction intervals become progressively tighter as the number of historical graphs increases. These findings indicate that the nonparametric s-RAM effectively captures the underlying structure governing the longest path outcomes.
@@
Figure 2. Prediction accuracy of nonparametric s-RAM on data generated from an underlying parametric s-RAM. Blue dots denote ground-truth values. Red ranges represent prediction intervals for the predicted probabilities by using nonparametric s-RAM.
6.2. Shortest Path Problem
Next, we evaluate the performance of s-RAM for the shortest path problem. Due to space limitations, we refer the reader to Section B.2.1 for details on the representability check, estimation, and prediction formulations, and to Section B.1 for a complete description of the data generation.
6.2.1. Explanatory ability of s-RAM in the shortest path problem
Generation of historical $\mathcal{X}^{(k)}$ and $\boldsymbol{p}^{(k)}$ instances. We follow the same experimental setup as in the longest path problem described in Section 6.1 to generate instances for evaluating the explanatory ability of s-RAM in the shortest path setting.
Representability test and estimation with $\boldsymbol{p}^{\mathcal{K}}$. The results of representability and estimation loss for the shortest path problem are reported in Table 3. Similar to the longest path case, Table 3 shows that s-RAM performs well across all tested scenarios despite model misspecification, and exhibits even stronger explanatory ability for the shortest path problem. In particular, s-RAM achieves high representability for $K\leq 15$. For example, when $K=5$, at least $90\%$ of the instances are s-RAM representable across all correlation structures, and the representable fraction remains above $70\%$ even when $K=15$. Moreover, the associated estimation error is substantially smaller: the average absolute deviation loss is consistently on the order of $10^{-4}$ or smaller, even in settings where exact representability is violated. This represents an order-of-magnitude improvement compared to the longest path problem. As in the longest path, representability decreases with $K$ due to the increasing number of feasibility constraints in (22) induced by additional historical observations. Overall, the framework remains robust across different correlation structures, with particularly stronger performance under independent and negatively correlated noise than positively correlated noise.
Table 3. Representation and estimation in the shortest path problem with Gaussian instances
K independent gaussian gaussian with -ve correlation gaussian with +ve correlation
fraction
of s-RAM
representable
instances
average absolute
deviation loss (se)
$(10^{-4})$
fraction
of s-RAM
representable
instances
average absolute
deviation loss (se)
$(10^{-4})$
fraction
of s-RAM
representable
instances
average absolute
deviation loss (se)
$(10^{-4})$
5 0.90 0.62 (0.53) 0.90 0.32 (0.23) 0.95 0.04 (0.04)
10 0.70 3.98 (2.25) 0.95 0.85 (0.83) 0.70 2.66 (1.38)
15 0.70 3.61 (2.12) 0.75 2.64 (2.09) 0.75 3.60 (1.93)
20 0.60 3.08 (1.32) 0.65 1.02 (0.44) 0.15 8.68 (3.40)
6.2.2. Predictive ability of s-RAM in the shortest path problem
We again evaluate the predictive accuracy of the nonparametric s-RAM on counterfactual tasks using data generated from parametric s-RAMs. For the shortest path problem, we focus on predicting the bottleneck congestion (the maximum flow on any link) on a new graph $\boldsymbol{G}^{\text{new}}$.
Generation of training instance $\boldsymbol{p}^{(k)}$ and ground truth in the shortest path. We adopt a data generation procedure that is largely identical to that used for the longest path problem, except that we vary the parametric noise distributions in order to test the robustness of the nonparametric s-RAM to different distributional assumptions. Specifically, for the shortest path problem, the marginals follow non-identical uniform distributions, where the lower bounds are sampled uniformly from $(1,3)$, and the upper bounds are sampled uniformly from $(5,10)$. All other edge costs are sampled independently and uniformly from $[3,5]$. Directed acyclic graphs are generated using the same procedure as in the longest path experiments. Then, we use formulation (19) to generate training instance $\boldsymbol{p}^{(k)}$. We evaluate the prediction performance of the nonparametric s-RAM across collection sizes $K\in\{5,10,20\}$. For each value of $K$, we generate $20$ training instances. Given a new graph $\boldsymbol{G}^{\text{new}}=(s\cup t\cup\boldsymbol{V}^{\text{new}},\boldsymbol{E}^{\text{new}})$, let $\boldsymbol{x}^{\text{true}}$ denote the ground truth flow for $\boldsymbol{G}^{\text{new}}$. To generate $\boldsymbol{x}^{\text{true}}$, we use the same parameter values as in the training instances together with formulation (19) to compute the true flow on the new graph. Then, the corresponding counterfactual ground-truth bottleneck congestion is then given by $\max_{e\in\boldsymbol{E}^{\text{new}}}x_{e}^{\text{true}}$.
Prediction with nonparametric s-RAM in the shortest path. For the shortest path problem, we use the s-RAM best-case estimate of the bottleneck congestion, defined as the minimum achievable maximum edge flow over all s-RAMs consistent with the observed data on the new graph. Formally, this quantity is given by $\min_{\boldsymbol{x}\in\mathcal{U}{\text{new}}}\max_{e\in\boldsymbol{E}^{\text{new}}}x_{e}$. Recall that $\mathcal{U}{\text{new}}$ denotes the feasible s-RAM nonparametric prediction region defined in Section 4.1. As shown in Figure 3, in the shortest path problem, the optimistic bottleneck congestion predicted by nonparametric s-RAM converges to the true bottleneck congestion as more graphs are observed.